OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -8 - 3 + 2 + 7 + ... + 67

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

this one is arithmetic, right? so what is the common difference?

OpenStudy (anonymous):

5

OpenStudy (irishboy123):

yes so \(a_1 = -8\) \(a_2 = -8 + 1(5)\) \(a_3 = -8 + 2(5)\) we will want a general term for \(a_n\), the nth term in this sequence

OpenStudy (irishboy123):

can you have a go at that?

OpenStudy (irishboy123):



OpenStudy (anonymous):

Would it be:\[\sum_{n=0}^{\infty}(-8+5n) \]

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

if we are starting at n = 1, you need a small tweak

OpenStudy (irishboy123):

if term 1 is \(a_1\), with \(n = 1\)

OpenStudy (anonymous):

sum_{n=0}^{15}(-8+5n) ?

OpenStudy (anonymous):

\[\sum_{n=0}^{15}(-8+5n)\]

OpenStudy (irishboy123):

\(a_1=−8 = -8 +5(1-1)\) \(a_2=−8+(5)(2-1)\) \(a_3=−8+(5)(3-1)\)

OpenStudy (irishboy123):

\(a_n = ??\)

OpenStudy (anonymous):

-8+5n

OpenStudy (irishboy123):

-8+5(n - ??)

OpenStudy (anonymous):

That's not an answer choice though. These are my answer choices: A. \[\sum_{n=0}^{15}(-8+5n)\] B. \[\sum_{n=0}^{\infty}(-40n)\] C.\[\sum_{n-0}^{15}(-40n)\] D. \[\sum_{n=0}^{\infty}(-8+5n)\]

OpenStudy (irishboy123):

OK, they're doing it that way, with the first term as \(a_0\) not \(a_1\) in which case you go with your suggestion, \(-8+5n\) and checking the value of \(n\) for the last term: \(-8+5n = 67 \implies n = 15\)

OpenStudy (anonymous):

Ok thanks! So it would be A?

OpenStudy (irishboy123):

\( \huge \checkmark \)